Solve the inequality

x-3 ≤ 2x+5 < -x+13

√(x2-2x-3) ≥ Ix-2I - 2x


1
Expert's answer
2021-08-31T13:47:26-0400

1.


"\begin{cases}
 x-3\leq2x+5 \\
 2x+5<-x+13
\end{cases}"

"\begin{cases}
 x\geq-8 \\
 3x<8
\end{cases}"

"\begin{cases}
 x\geq-8 \\
 x<8/3
\end{cases}"

Answer: "-8\leq x<8/3"


"x\in[-8, 8/3)"

2.


"\sqrt{x^2-2x-3} ≥ |x-2|- 2x"


"x^2-2x-3\geq0=>(x+1)(x-3)\geq0"

"x\leq-1\ or\ x\geq3"

"x\leq-1"


"\sqrt{x^2-2x-3} ≥ 2-x-2x"

"\sqrt{x^2-2x-3} ≥ 2-3x"

"2-3x>0, if\ x\leq-1"

"x^2-2x-3\geq(2-3x)^2"

"x^2-2x-3\geq4-12x+9x^2"


"8x^2-10x+7\leq0"

"D=(-10)^2-4(8)(7)=-124<0"

"8>0, D<0, 8x^2-10x+7>0, x\in \R"

There are no solutions for "x\leq-1."


"x\geq3"



"\sqrt{x^2-2x-3} ≥x-2-2x"

"\sqrt{x^2-2x-3} ≥ -x-2"

"-x-2<0, if\ x\geq3"

Then

"\sqrt{x^2-2x-3}\geq0>-x-2, x\geq3"

Answer: "x\geq3"


"x\in[3,\infin)"